ON DIFFUSION-INDUCED BLOWUPS IN A COOPERATIVE MODEL

YUAN LOU WEI-MING NI

Abstract. The asymptotic behaviors of solutions to a cooperative model with diffusion are studied. Sufficient conditions are given for solutions to blow up. In particular, we show that with the help of diffusion, the population of two cooperating species may blow up in finite time, although both species with the same initial data could die out at every point of their habitat if neither of them migrates.